7:56 AM Tue Nov 13 15. Your vial of Potassium Chloride is labeled 5 mEq/5 ml. Your patient must receive Potassium Chloride 10mEq, by mouth, three times a day. You would give dose? tsp per 16. You have a 500 ml bottle of dexamethasone. The label reads 0.5 mg per 5 ml. In order to administer dexamethasone 4 mg po, you would give ml. 17. Your 50 ml bottle of medication is labeled: Sulfasoxazole susp. 250 mg/5 ml. You must administer 300 mg by mouth four times per day. each dose? How many ml would you administer for 300my 250 m 18. You have Ergonovine 325mg tablets on hand. 6. 8. 10, The doctors order reads to give Ergonovine 650mg twice daily x 2 days. How many tablets would you give? I TAS 325m3 19. The pharmacy sent a 2.5ml vial of streptomycin for your patient. There is a total of 1 gm streptomycin in the vial. You must give 800mg of streptomycin IM daily to your patient. How many ml would you administer for the dose?
Answer #1

Q.15. Vial labelled =5mEq/5ml

patient requires 10mEq by mouth three times a day

the required tsp perdose?

Single dose is 10mEq means 10 ml

we know that 1 tsp=4.928 ml means almost 5ml

Hence the required tsp perdose will be 10/5=2tsp

Q.16. Available 500 ml bottle of dexamethasone

0.5mg/5ml

required dose 4mg po

how many ml should take?

0.5mg in 5ml

1mg in 10 ml

hence 4mg will be in 40ml

Q.17. available 250mg/5ml

required 300mg by mouth qd

how many ml per dose ?

250mg/5ml

means50mg/ml

required is 300mg per dose hence requiredml will be 300/50=6ml

Q.18. ergotamine 325 mg tablet is available, required 650mg twice daily for 2 days required tabs?

means 2 tabsX2=4 per day

hence for 2 days required tabs will be 4x2=8 tabs total

Q.19. in 2.5 ml vial of streptomycin there is 1 gm of streptomycin

required dose 800mg IM daily

how may ml should administer?

1000mg in 2.5ml

1mg will be in 2.5/1000

hence 800mg =(2.5/1000) x800

= 2ml

hence 2ml should adminester